# Impact of Calibration and Position Errors on Astrophysical Parameters of the Hi 21cm Signal

### v1

@@ 5 Emulating Hi 21cm Power Spectrum
The objective of developing ANN-based emulators for this work is to study the power spectrum, as these emulators can generate efficient and reliable EoR models. These models can then be utilized as substitutes for computationally intensive simulations in Bayesian parameter inference. In addition to this, a non-parametric feature of ANN emulators is the ability to replicate various signal features without relying on their specific parametric characteristics, as they are solely dependent on the training data sets. This capability allows us to utilize the features directly without the need to remove or apply any cleaning algorithms, thereby enabling the inference of associated label parameters. In this study, we are developing emulators by training with different PS observed by different interferometric arrays like SKA Low and HERA. We also train the emulator for the theoretical power spectrum, which one can say is the power spectrum for the perfect array interferometric condition. We use the ANN model to build these emulators. To develop ANN architecture, we are using the Python-based deep learning Keras API, and standard sci-kit learn [101]. To achieve significant accurate PS from the emulator for test set of EoR parameter we have to train the emulator with the sufficient number of the training datasets. To construct an optimal training data set, we sample the parameter space with Latin Hypercube Sampling. Latin Hypercube sampling method samples the multi-dimensional parameter space such that no two parameters share the same value in the parameter space, providing an all-unique set of parameters. We sampled the EoR parameters range into 300 grid points using Latin Hypercube sampling to build the training and test datasets. These 300 sample data points given as input to construct the lightcones details are described in section 2.1, which is further used to construct the observed power spectrum using the 21cm E2E pipeline. To develop construct training and test datasets for the theoretical PS, we use 21cmFast. The parameters range we follow to construct the training datasets are = (10 Mpc, 60 Mpc ), = (4.5 K, 6.0 K ), and = (10, 60). We construct training and test data sets of observed/theoretical power spectrum () by following these sample parameters for 6 different modes. We use 270 sample data sets to train and validate the ANN model and 30 data sets to test the ANN emulator. We constructed the network architecture by following Python-based Keras’ Sequential API. The network consists of an input layer consisting 3 neurons matching with the training 3 EoR key parameter (, , ) and two hidden layers with 28 and 14 neurons, respectively, each activated by the ’elu’ activation function. The output layer has 6 neurons to predict the observed PS. The input PS data sets are normalized using the’ StandardScaler function, and corresponding EoR parameters are normalized using MinMaxScaler, which is available in sklearn. We trained this ANN model using different power spectra (PS) observed by various interferometric arrays, such as SKA Low and HERA, to create individual emulators for each array. Fig. 3 presents a comparison between the true test set 21-cm PS, shown as solid lines, and the emulated PS by the ANN, represented by dots. For this comparison, we randomly selected five test sets of power spectra and compared them with the predicted PS from the emulators in each case.
@@ 6 Bayesian Inference of EoR Parameters
We design our pipeline to extract the most likely EoR parameter values using MCMC for the given log-likelihood. We used the CosmoHammer [102] Python package, utilizing an affine-invariant [103] Markov Chain Monte Carlo (MCMC) ensemble sampler, to conduct model parameter estimation. To test the performance of the pipeline we input either the binned theoretical 21-cm power spectrum or simulated observed power spectrum.
@@ 7 Error Covariances of Power Spectrum
The error estimate in the observables is essential to predict the posterior of the parameters. The EoR 21cm signal has varoius contamination from the foreground, systematics, noise, calibration errors etc. To estimate the error covariance of the power spectrum (PS), we assume that the foregrounds have been perfectly modeled and entirely removed from the observed signal. The remaining observed signal consists of the EoR 21cm signal and Gaussian thermal noise. Hence, in this analysis, the total covariance () is defined as the sum of the sample variance () and the thermal noise variance ().
@@ 8 Result
This section describes simulation outcomes and discusses their relevance to actual observations. The results for each considered configuration are detailed in individual sections below.
@@ 8.1 Perfect Observing Condition
In this study, we derived astrophysical characteristics from various interferometric arrays’ observed power spectra (PS). We assumed that foregrounds were perfectly removed from the observed PS and that there were no calibration or positional errors, as shown in Fig. 2. To achieve this, we simulated observations using the 21cm E2E pipeline with only input Hi lightcone maps and calculated the PS from the resulting observed Hi visibility. To derive the astrophysical parameters, we employed an emulator-based Markov Chain Monte Carlo (MCMC) pipeline. The advantage of employing this emulator-based MCMC pipeline lies in its ability to reduce computational time significantly. Furthermore, as the emulators are trained on observed power spectra that inherently incorporate telescope effects, they allow for direct parameter inference without the need for separate corrections for these effects. The inferred astrophysical parameters for various interferometric arrays are depicted in Fig.5. For all cases, we found that the two parameters, and , closely matched the actual values. However, the third parameter, , was not correctly constrained in any of the cases. To investigate this issue, we examined whether the lack of constraint could be due to the effects introduced by the telescope. Additionally, we compared these results with the inferred astrophysical parameters derived from theoretical power spectra. A similar trend was observed in this case as well. Specifically, and closely matched the actual values, as shown in Fig.4. However, remained unconstrained in all scenarios, which may be attributed to the degeneracy of the parameter. To further assess the robustness of this pipeline, we inferred parameters for two additional power spectra generated using astrophysical parameters near the boundary values of those used in the training dataset. The inferred parameters for these power spectra are presented in Fig.7, with set 2 on the left and set 3 on the right, as detailed in Appendix A. Similar to the initial case, the astrophysical parameters and were well-constrained and closely aligned with the true values in most instances, as depicted in Fig.7. However, the parameter remained poorly constrained across both sets.
@@
We also noticed that the power spectrum from set 1, whose astrophysical parameters are around the centre of the boundaries, generated more accurate conclusions across all interferometric measured PS. By contrast, the power spectra generated in close proximity to the boundary values of the astrophysical parameters yielded inferred parameter mean that differed from the actual value. This mismatch could be due to the ANN emulator being trained on restricted datasets. Potential future development of a more precise emulation could involve training it with a more extensive dataset.
@@ 8.2 Imperfect Observing Conditions
Sensitive radio observations like the ones targeting the Hi signal from CD/EoR is prone to contamination from a large number of sources. Small errors in initial processing steps like calibration can propagate into the final power spectrum estimates, resulting in misinterpretation of the obtained results. In [1], the authors studied the effects of calibration errors and positions errors in the recovery of the 21-cm power spectrum. In this study, we follow up by exploring the effect of such errors on the recovery of astrophysical parameters. The visual comparison of the PS was previously conducted by [1], where they examined both 2D and 1D residual PS with gain errors and position errors. They also investigated the variation of root mean square (RMS) in the residual image across different percentages of calibration errors and position errors, comparing it with the signal level and thermal noise. In the previous study by [1], it was found that a gain error threshold of 0.01% leads to a high image RMS for both the SKA-Low and HERA interferometric arrays, potentially causing confusion or even obscuring the detection of faint cosmological signals. Similarly, they explored the variation of RMS in the residual image across different position errors, comparing it with the signal level and thermal noise. In their study, [1] found that a position error threshold of 0.48” leads to a high image RMS, potentially causing confusion or even obscuring the detection of faint cosmological signals. In our study, we compared the corrupted 1D power spectrum (PS) with the original PS and estimated the astrophysical parameters from the corrupted PS for both scenarios: when it was corrupted by gain calibration errors and when it was corrupted by position errors.
@@ 8.2.1 Gain Calibration Errors
Gain calibration errors were incorporated following equation 1 from [1], applied per antenna rather than per baseline, affecting the entire k range.
@@
SKA Low: To assess the influence of calibration errors on the observed power spectrum (PS) and their subsequent impact on associated astrophysical parameters for SKA-Low, we introduced various gain errors as detailed in [1]. The resulting residual PS was visually compared with the actual PS, depicted in the top left panel of Fig.5. The parameter obtained from the PS analysis with SKA-Low, incorporating actual signal and PS for gain error residuals, is depicted in the top right panel of Fig. 5. Upon examining the plotted residual PS spectrum, we observed that for gain errors of 0.001 %, the residual PS overlaps with the true PS. For residual PS with gain errors of 0.01%, we observe overlapping at lower k modes, but with minor deviation at higher k modes. However, when deriving astrophysical parameters from the residual power spectrum, we found that for a gain error of 0.001%, the derived parameters match those derived from the true PS. In contrast, for PS with the gain error of 0.01%, the derived parameters exhibit significant differences from the actual values. This indicates that astrophysical parameters are highly sensitive to gain errors, and even minor deviations in the PS due to gain errors can severely impact the estimation of parameters, introducing bias. The sensitivity of derived parameters to gain errors clearly suggests that the gain error threshold should be much lower than 0.01%, ideally closer to 0.001%. If the gain error exceeds 0.001%, it is advisable to model them to prevent bias in the derived parameter values.
HERA: HERA, a crucial instrument for studying the Epoch of Reionization (EoR), has begun its observations and recently produced its first upper limit based on initial phase data [27]. To assess the impact of calibration errors on the observed power spectrum (PS) and the corresponding astrophysical parameters for HERA, we introduced various gain errors similar to those observed in SKA-Low. The residual power spectrum (PS) due to different gain errors is presented in the bottom left panels of Fig. 5, while the posterior distributions of model parameters derived from these PS measurements are shown in the bottom right panels of Fig. 5. Upon examining the residual PS spectrum, we found that when the gain error is 0.001%, the residual PS closely aligns with the true PS. But when the gain error increases to 0.01%, the residual PS overlaps with the true PS at lower k modes but diverges slightly at higher k modes.
The derived astrophysical parameters are consistent with the true PS for a gain error of 0.001%, but significant discrepancies emerge at a 0.01% gain error. This indicates that even minor deviations in the PS due to gain errors can considerably affect parameter estimation, leading to bias. Thus, it is essential for gain error thresholds to remain below 0.01%, ideally closer to 0.001%. If gain errors exceed 0.001%, they should be carefully modeled to avoid bias in the parameter values.
8.2.2 Position Error
@@
To evaluate the impact of position errors in the global sky model, simulations were conducted using inaccurate sky models as described in [1]. Position errors were modeled with zero-mean Gaussian distributions and varying standard deviations (up to 0.048”, 0.48”, and 4.8”) and applied to the right ascension (RA) of sources, resulting in a new catalog with positional inaccuracies. Residuals were then obtained following equation 4 from [1], by subtracting the corrupted sky from the true sky, and these residuals were used to analyze the effects on the PS.
@@
SKA Low: Similar to calibration errors, we evaluated the impact of position errors on the observed power spectrum (PS) and their effect on the inferred astrophysical parameters for SKA-Low. Different position errors, as described above, were introduced, and the resulting residual PS was compared with the true PS, as shown in the top-left panel of Fig. 6. The parameters obtained from the PS, which incorporates actual signals and position error residuals, are depicted in the top right panel of Fig. 6. Upon examining the plotted residual PS spectrum, we found that for a position error of 0.048”, the residual PS overlaps with the true PS. At lower k modes, the residual PS with this position error overlaps well with the true PS, but at higher k modes, there is a minor deviation. This occurs because real-space scales are more affected by arcsecond-level errors at higher k values. However, since we used the entire k range to derive astrophysical parameters from the residual power spectrum, we observed that for a position error of 0.048”, the derived parameters closely match those obtained from the true PS. Similarly, for a position error of 0.48”, the derived parameters exhibit substantial deviations from the true values. This indicates that astrophysical parameters are highly sensitive to position errors, and even small deviations in the PS due to these errors can severely impact parameter estimation, introducing bias. The sensitivity of derived parameters to position errors, similar to gain errors, suggests that the acceptable position error threshold should be much lower than 0.48”, ideally around 0.048”. If the position error exceeds 0.048”, modeling the errors becomes crucial to prevent bias in the derived parameter values.
HERA : To assess the impact of position errors on the observed power spectrum (PS) and their influence on astrophysical parameters for HERA, we introduced various position errors similar to those in SKA-Low. The residual PS was visually compared with the actual PS, as shown in the bottom-left panel of Fig.6. The parameters obtained from the PS analysis, incorporating actual signal and PS for position error residuals, are depicted in the bottom-right panel of Fig.6. We observed that for position errors of 0.048”, the residual PS overlaps with the true PS, with overlap at lower k modes and very minor deviations at higher k modes. However, for a position error of 0.48”, the derived parameters significantly differed from the actual values, indicating high sensitivity to position errors. This suggests that the position error threshold should be much lower than 0.48”, ideally closer to 0.048”. If position errors exceed 0.048”, they should be modeled to prevent bias in parameter estimation.

### v2

@@ 5 Emulating Hi 21cm Power Spectrum
The objective of developing ANN-based emulators for this work is to study the power spectrum, as these emulators can generate efficient and reliable EoR models. ANNs are a class of machine learning models inspired by the neural architecture of the human brain. They consist of multiple layers of interconnected computational units, referred to as neurons, which process and propagate information through the network. Each neuron applies a transformation to its input via a mathematical function known as an activation function, and passes the result to subsequent layers. Hidden layers those situated between the input and output layers facilitate the learning of complex and non-linear relationships within the data. Commonly used activation functions such as the Rectified Linear Unit (ReLU) and the Exponential Linear Unit (ELU) introduce the necessary non-linearity for the network to approximate highly intricate mappings between inputs and outputs.
@@
These models can then be utilized as substitutes for computationally intensive simulations in Bayesian parameter inference. In addition to this, a non-parametric feature of ANN emulators is the ability to replicate various signal features without relying on their specific parametric characteristics, as they are solely dependent on the training data sets. This capability allows us to utilize the features directly without the need to remove or apply any cleaning algorithms, thereby enabling the inference of associated label parameters. In this study, we are developing an emulator by training it on a set of mock observed power spectra generated from simulated observations with the SKA-Low interferometric array. We also train the emulator on the theoretical power spectrum, that is, in the absence of both instrumental effects and array configuration influences. We use the ANN model to build these emulators. To develop ANN architecture, we are using the Python-based deep learning Keras API, and standard sci-kit learn [101]. To achieve significant accurate PS from the emulator for test set of EoR parameter we have to train the emulator with the sufficient number of the training datasets. To construct an optimal training data set, we sample the parameter space with Latin Hypercube Sampling. Latin Hypercube sampling method samples the multi-dimensional parameter space such that no two parameters share the same value in the parameter space, providing an all-unique set of parameters. We generated 300 unique combinations of astrophysical parameters using Latin Hypercube Sampling from the defined parameter ranges to construct the training and test datasets. These sampled parameter sets were used to simulate lightcones, as detailed in Section 2.1, and subsequently to compute the observed power spectra using the 21cm end-to-end (E2E) pipeline. To develop construct training and test datasets for the theoretical PS, we use 21cmFast. The parameters range we follow to construct the training datasets are = (10 Mpc, 60 Mpc ), = (4.5 K, 6.0 K ), and = (10, 60). We construct training and test data sets of observed/theoretical power spectrum () by following these sample parameters for 6 different modes. We use 270 PS data sets to train and validate the ANN model and 30 data sets to test the ANN emulator. We constructed the network architecture by following Python-based Keras’ Sequential API. The network consists of an input layer consisting 3 neurons matching with the training 3 EoR key parameter (, , ) and two hidden layers with 28 and 14 neurons, respectively, each activated by the ’elu’ activation function. The number of neurons in the hidden layers is determined empirically through trial and error, based on the configuration that yields the best performance in terms of accuracy and robustness. The output layer has 6 neurons to predict the observed PS. To ensure effective model training and convergence, the input PS values are standardized using the StandardScaler function, which removes the mean and scales the features to unit variance. The corresponding EoR parameters are normalized to the [0, 1] range using MinMaxScaler. Both scalers are implemented using the sklearn library. This preprocessing step is essential for bringing all input features to a comparable scale, thereby enhancing training stability and learning efficiency. We trained this ANN model using different PS observed by SKA-Low interferometric arrays, to create emulator. Fig. 3 presents a comparison between the true test set 21-cm PS, shown as solid lines, and the emulated PS by the ANN, represented by dots. For this comparison, we randomly selected five test sets of power spectra and compared them with the predicted PS from the emulators in each case.
@@ 6 Bayesian Inference of EoR Parameters
We design our pipeline to extract the most likely EoR parameter values using MCMC for the given log-likelihood. We used the CosmoHammer [102] Python package, utilizing an affine-invariant [103] Markov Chain Monte Carlo (MCMC) ensemble sampler, to conduct model parameter estimation. We have tested the performance of the pipeline using both the theoretical power spectrum and the simulated observed power spectrum. Further details are discussed in Section 8.
@@ 7 Error Covariances of Power Spectrum
The error estimate in the observables is essential to predict the posterior of the parameters. The EoR 21cm signal has varoius contamination from the foreground, systematics, noise, calibration errors etc. To estimate the error covariance of the power spectrum (PS), we assume that the foregrounds have been perfectly modeled and entirely removed from the observed signal. This perfect modelling assume of the sky and instrument to establish a baseline for evaluating performance under idealized conditions. This allows us to isolate the intrinsic capabilities and limitations of the methodology without complications from model inaccuracies. In the following section, we introduce modelling imperfections such as gain calibration and sky position errors to assess the robustness of the method and quantify their impact on power spectrum estimation. This comparison highlights the potential biases that arise when instrumental and modelling errors are neglected, motivating future work on mitigation strategies. Under the perfect modelling assumption, the remaining observed signal consists of the EoR 21cm signal and Gaussian thermal noise. Hence, in this analysis, the total covariance () is defined as the sum of the sample variance () and the thermal noise variance ().
@@ 8 Result
This section presents the simulation results and discusses their implications for real observations. Detailed outcomes for each case are provided in the following subsections.
@@ 8.1 Perfect Observing Condition
In this study, we infer astrophysical parameters from the power spectra (PS) obtained using the SKA-Low interferometric array. We assumed that foregrounds were perfectly removed from the observed PS and that there were no calibration or positional errors, as shown in Fig. 2. To achieve this, we simulated observations using the 21cm E2E pipeline with only input Hi lightcone maps and calculated the PS from the resulting observed Hi visibility. To derive the astrophysical parameters, we employed an emulator-based Markov Chain Monte Carlo (MCMC) pipeline. The advantage of employing this emulator-based MCMC pipeline lies in its ability to reduce computational time significantly. Furthermore, since the emulators are trained on observed power spectra that inherently incorporate the effects of the telescope layout, they enable direct parameter inference without requiring separate corrections for these effects. The astrophysical parameters inferred using the SKA-Low interferometric array are presented in Fig. 4. In this case, we found that the two parameters, and , closely matched the actual values. However, the third parameter, , was not correctly constrained. To investigate this issue, we examined whether the lack of constraint could be due to the effects introduced by the telescope layout. Additionally, we compared this result with the inferred astrophysical parameters derived from theoretical power spectra. A similar trend was observed in this case as well. Specifically, and closely matched the actual values, as shown in Fig. 4. However, remained unconstrained, which may be attributed to the degeneracy of the parameter. To further assess the robustness of this pipeline, we inferred parameters for two additional power spectra generated using astrophysical parameters near the boundary values of those used in the training dataset. The inferred parameters for these power spectra are presented in Fig. 7, with set 2 on the left and set 3 on the right, as detailed in Appendix A. Similar to the initial case, the astrophysical parameters and were well-constrained and closely aligned with the true values in most instances, as depicted in Fig. 7. However, the parameter remained poorly constrained across both sets. As noted in [106], this degeneracy arises around redshift , where a population of low-mass galaxies with brighter stellar populations can replicate the fiducial observation. These galaxies tend to ionize the IGM earlier, and as a result, combining observations across multiple redshifts and incorporating the evolution of the ionization process can help break this degeneracy. In future work, we plan to test this explicitly by including multi-redshift observations to assess the improvement in parameter recovery.
@@
We also noticed that the power spectrum from set 1, whose astrophysical parameters are around the centre of the boundaries, generated more accurate conclusions. By contrast, the power spectra generated in close proximity to the boundary values of the astrophysical parameters yielded inferred parameter mean that slightly differed from the actual value. For example, in Set 3, the true value of is 57.42, while the inferred mean using the theoretical power spectrum is 50.8. Using the SKA power spectrum, the inferred mean is 59.8. Although these values differ from the true mean, they still lie within of the inferred uncertainty range (see Table 2). This mismatch could be due to the ANN emulator being trained on less training datasets. Potential future development of a more precise emulation could involve training it with a more extensive dataset that cover the parameter space in the effective manner.
@@ 8.2 Imperfect Observing Conditions
Sensitive radio observations targeting the Hi signal from the CD and EoR are highly susceptible to contamination from various sources. Small errors in early data processing steps, such as calibration, can propagate into the final power spectrum (PS) estimates, potentially leading to misinterpretation of the results. In [1], the authors investigated the effects of gain calibration errors and sky model position errors on the recovery of the 21-cm power spectrum. In this study, we build upon that work by examining how such errors impact the recovery of astrophysical parameters. [1] previously conducted a visual comparison of the power spectrum, analyzing both 2D and 1D residual PS in the presence of gain calibration and sky position errors. They also evaluated how the root mean square (RMS) of the residual image varies with different levels of gain calibration and position errors, comparing these values against the signal and thermal noise levels. Their findings showed that a gain calibration error as small as 0.01% could result in a high image RMS in SKA-Low observations, potentially causing confusion or even masking faint cosmological signals. Likewise, a sky model position error exceeding 0.48 arcseconds was found to significantly elevate the image RMS, leading to similar challenges in signal detection. In our analysis, we extend this investigation by comparing the corrupted 1D power spectrum with the original and estimating the astrophysical parameters from the corrupted spectra in two scenarios: one with gain calibration errors and another with sky model position errors.
@@ 8.2.1 Gain Calibration Errors
Gain calibration errors were incorporated by applying complex gain perturbations to each antenna individually, as expressed in Equation 8.1, following the methodology of [1]. These per-antenna errors propagate through the visibilities since each visibility is formed from a pair of antennas and thus impact the entire space range. The visibility measured between a pair of antennas and , denoted as , is given by:
Here, is the true visibility from the sky, while and are the complex gains of antennas and , respectively. Although gain calibration attempts to estimate and correct for these instrumental effects, it is typically imperfect, leaving behind residual errors that can propagate through the analysis and bias the recovered signal.
The complex gain for each antenna can be modeled as:
where is the nominal amplitude and the phase (in radians), with and representing the residual errors in amplitude and phase, respectively. In an ideal scenario, and , yielding a perfect gain of 1. However, with calibration imperfections, the gain simplifies to:
@@
To assess the influence of gain calibration errors on the observed power spectrum (PS) and their subsequent impact on associated astrophysical parameters for SKA-Low, we introduced various gain calibration errors as detailed in [1]. The resulting residual PS was visually compared with the actual PS, depicted in the top left panel of Fig. 5. The parameter obtained from the PS analysis with SKA-Low, incorporating actual signal and PS for gain error residuals, is depicted in the top right panel of Fig. 5. Upon examining the plotted residual PS spectrum, we observed that for gain calibration error of 0.001 %, the residual PS overlaps with the true PS. For residual PS with a gain calibration error of 0.01%, we observe overlapping at lower k modes but with minor deviation at higher k modes. However, when deriving astrophysical parameters from the residual power spectrum, we found that for a gain calibration error of 0.001%, the derived parameters match those derived from the true PS. In contrast, for PS with the gain calibration error of 0.01%, the derived parameters exhibit significant differences from the actual values. This indicates that astrophysical parameters are highly sensitive to gain calibration errors, and even minor deviations in the PS due to gain calibration errors can severely impact the estimation of parameters, introducing bias. The sensitivity of derived parameters to gain calibration errors clearly suggests that the gain calibration error threshold should be much lower than 0.01%, ideally closer to 0.001%. If the gain calibration error exceeds 0.001%, it is advisable to model them to prevent bias in the derived parameter values.
8.2.2 Sky Model Position Error
@@
To evaluate the impact of sky model position errors in the sky model, simulations were conducted using inaccurate sky models as described in [1]. Sky model position errors were modeled with zero-mean Gaussian distributions and varying standard deviations (up to 0.048”, 0.48”, and 4.8”) and applied to the right ascension (RA) of sources, resulting in a new catalog with positional inaccuracies. Residuals were then obtained following equation 4 from [1], by subtracting the corrupted sky from the true sky, and these residuals were used to analyze the effects on the PS. Unlike gain calibration errors which arise from instrumental calibration inaccuracies sky model position errors are rooted in prior assumptions about the sky and do not depend on the calibration solution itself. However, they can still introduce significant foreground residuals that mimic spectral structure, making them especially relevant for 21cm cosmology. This distinction is important, as position errors affect subtraction fidelity rather than the correction of instrumental response, and thus represent a separate and critical source of systematic error.
@@
Similar to gain calibration errors, we evaluated the impact of sky model position errors on the observed power spectrum (PS) and their effect on the inferred astrophysical parameters for SKA-Low. Different sky model position errors, as described above, were introduced, and the resulting residual PS was compared with the true PS, as shown in the top-left panel of Fig. 6. The parameters obtained from the PS, which incorporates actual signals and position error residuals, are depicted in the top right panel of Fig. 6. Upon examining the plotted residual PS spectrum, we found that for a sky model position error of 0.048”, the residual PS overlaps with the true PS. At lower k modes, the residual PS with this position error overlaps well with the true PS, but at higher k modes, there is a minor deviation. This occurs because real-space scales are more affected by arcsecond-level errors at higher k values. However, since we used the entire k range to derive astrophysical parameters from the residual power spectrum, we observed that for a position error of 0.048”, the derived parameters closely match those obtained from the true PS. Similarly, for a sky model position error of 0.48”, the derived parameters exhibit substantial deviations from the true values. This indicates that astrophysical parameters are highly sensitive to sky model position errors, and even small deviations in the PS due to these errors can severely impact parameter estimation, introducing bias. The sensitivity of derived parameters to sky model position errors, similar to gain calibration errors, suggests that the acceptable sky model position error threshold should be much lower than 0.48”, ideally around 0.048”. If the sky model position error exceeds 0.048”, modeling the errors becomes crucial to prevent bias in the derived parameter values.